OpenStudy (elise_a18):

I'm about to need tons of help with checking answers. If your willing please help me. There will be lots, this is Algebra I level btw

OpenStudy (anonymous):

I can help but I suck at it :)

OpenStudy (elise_a18):

okay thanks at least. I'll tag you in them once i pick out my answers :)

OpenStudy (anonymous):

ok hav fun :)

OpenStudy (elise_a18):

recursive function is shown below: f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 4, -4, -12, -20, -28, ... 4, 4, 12, 20, 28, ... 4, 12, 20, 28, 36, ... 4, -12, -20, -28, -36, ... is it C?

OpenStudy (elise_a18):

@Godlovesme @samanthagreer

OpenStudy (anonymous):

I hav 2 go but its c

OpenStudy (elise_a18):

oh kk

OpenStudy (elise_a18):

@AllTehMaffs

OpenStudy (anonymous):

So you don't even have to look at any functions other than f(2)! If we're looking at f(2) (with n=2), what's f(n-1)?

OpenStudy (elise_a18):

f(1)

OpenStudy (anonymous):

Yup yup. And we have a definition of f(1). So if n=2, and if \[ f(n) = f(n-1) -8\], then we have \[f(2) = f(1) - 8\] and what does that yield?

OpenStudy (anonymous):

(it's not C)

OpenStudy (elise_a18):

do it's be f(-7) but none of them start wih that :c

OpenStudy (anonymous):

Close! (and a preemptive apology for the long response) So the notation \[ f(n) \] doesn't mean the same thing as f*n It means "f of n." "f of n" means "what the function is when you plug in n = some number." For \[ f(1)\] the problem says that \[f(1) = 4\] So! If we're looking at \[f(2)\] it means the same thing saying \[f(n=2)\] So you have to plug everything back into the defined function - aka, the defined pattern (recursive functions start with a value then do something to that the make a new pattern). So! For n=2 \[f(n) = f(n-1)-8\] means \[f(2) = f(2-1) - 8\] \[=f(2) = f(1)-8\] and we already know that \[f(1) = 4\] If you replace \[f(1)\] with \[4\] then what does that look like?

OpenStudy (elise_a18):

Oh so it's A?

OpenStudy (elise_a18):

no wait

OpenStudy (elise_a18):

D

OpenStudy (elise_a18):

DX gah i dont know

OpenStudy (anonymous):

^_^ Why did you say A the first time?

OpenStudy (elise_a18):

Mm, because it looked right for some reason ;n;

OpenStudy (elise_a18):

>n<

OpenStudy (anonymous):

Trust your gut! If \[f(1) = 4\] and \[f(n) = f(n-1) - 8\] then \[f(2) = f(2-1)-8\] so \[f(2) = f(1) - 8\] and what does \[f(1)=?\] (it's silly and roundabout, but we're dealing with recursive functions so it's okay ^_~)

OpenStudy (elise_a18):

you gave it to me i think at the top f(1)=4

OpenStudy (elise_a18):

and since it's positive you'd be increasing my 4?

OpenStudy (elise_a18):

-3?

OpenStudy (elise_a18):

i really have no idea how these functions work

OpenStudy (elise_a18):

I've forgotten ;n;

OpenStudy (anonymous):

That doesn't mean you're gonna get away with not trying ^_^ And the best part about math is that everything is pretty similar, so even if you've forgotten on part the rest theoretically falls into place if you dwell long enough. If \[f(1) = 4\] then \[f(2) = f(1)-8\] means \[f(2) = 4 - 8\]

OpenStudy (elise_a18):

Oh I See it! so you'd just replace f(1) with the equivalent?

OpenStudy (anonymous):

Yup yup yup ^_^

OpenStudy (elise_a18):

i son't see how we got f(2) though

OpenStudy (anonymous):

It's another "equivalent." if you look at it as \[y = x-8\] and \[x=4\] what does that make y?

OpenStudy (anonymous):

If \[f(n) = f(n-1) - 8\] then we're going to say first that \[n=2\] Then everywhere you see an "n" you replace it with a 2, which gives \[f(2) = f(2-1)-8\] Remenber that \[f(1)=4\] is a definition for the problem, so everywhere you see the symbol \[f(1)\] you can substitute in a 4. That gives us \[f(2) = f(2-1)- 8 \longrightarrow f(2) = f(1) -8 \longrightarrow f(2) = 4-8 \longrightaarow f(2) = -4\] We can show that the answer is correct by keeping the pattern going. Just like \[f(1) = 4\] We've shown that \[f(2) = -4\] So if we look back at the recursive equation \[f(n) = f(n-1)-8\] We're going to say that the next item in the sequence is n=3, giving \[f(3) = f(3-1) -8\] so \[f(3) = f(2)-8\] and we've shown that \[f(2) = -4\] so \[f(3) = (-4)-8\] meaning \[f(3) = -12\] and the first three terms of our recursive function are \[4, -4, -12, ...\] which only matches on of the options ^_^

OpenStudy (anonymous):

grrrrrr, sorry the weird typo should say \[f(2) = f(2-1) - 8 \longrightarrow f(2) = f(1) - 8 \longrightarrow f(2) = 4-8 \longrightarrow f(2) = -4\]

OpenStudy (anonymous):

Does that make sense, or is it still a little iffy?

OpenStudy (elise_a18):

Yep I understand it enough to solve other ones. Thanks :)